OpenStudy (anonymous):

A fireworks company has two types of rockets called Zinger 1 and Zinger 2. The polynomial -16t2 + 150t gives the height in feet of Zinger 1 at t seconds after launch. The polynomial -16t2 + 165t gives the height of Zinger 2 at t seconds after launch. If the rockets are launched at the same time and both explode 6 seconds after launch, how much higher is Zinger 2 than Zinger 1 when they explode?

OpenStudy (anonymous):

t=6 Zinger 1: -16 * t^2 + 150 * t = -16 * 6^2 + 150 * 6 =324 Zinger 2: -16 * t^2 + 165 * t = -16 * 6^2 + 165 * 6 = 414 414-324=90

OpenStudy (anonymous):

Oh, I see Thank you!!

OpenStudy (anonymous):

You're welcome :)